OpenStudy (anonymous):

find the equation of the line useing the following info show all your work (-3,-4).m=3

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

gradient or slope is 3 and the line goes through the point (-3,-4). Use the formula y-y1 = m(x-x1)

OpenStudy (jocelyn):

-3 is x-value and -4 is y-value. m=3 which is the slope. the slope-intercept form is y=mx + b. -4=3(-3) + b. -4= -9 + b +9 +9 b= 5 Equation: y= 3x + 5

OpenStudy (anonymous):

Ok It is a line so has an equation y = mx + c (m is the gradient c is the constant at which the line passes through the y axis on its graph) So we now use the coordinates for our x and y values along with the gradient. -4 = 3(-3) + c therefore -4 = -9 + c Adding 9 to both sides we get 5 = c So the equation of the line is y = 3x +5

OpenStudy (anonymous):

thanks